In this question, you should state clearly all the distributions that you use, together with the values of the appropriate parameters. A cup of cheese tea is prepared with two ingredients, cream cheese foam and tea. The barista preparing the cheese tea is not allowed to remove cream cheese foam nor tea once they have been poured into the cup. The amounts, in ml, of cream cheese foam and tea in one cup of cheese tea made by an experienced barista have independent normal distributions with means and standard deviations as shown in the table.
Mean amount (ml)Standard deviation (ml)
Cream cheese foam725
Tea4214

  1. (i)

    Given that the amount of cheese tea in each cup is expected to be at least 490 ml, find the probability that a randomly chosen cup of cheese tea made by the experienced barista meets this expectation.

  2. (ii)

    A cup of cheese tea is considered well-made if the amount of cream cheese foam is between 65 ml and 75 ml and the amount of tea is between 416 ml and 424 ml.

    Find the probability that, for ten randomly chosen cups of cheese tea made by the experienced barista, more than five cups are well-made.

  3. (iii)

    The cost price of cream cheese foam and tea is $20 per litre and $1.20 per litre respectively. A cup of cheese tea is sold at $4.90.

    Without factoring in other costs incurred in producing the cheese tea, find the probability that the profit for 100 randomly chosen cups of cheese tea made by the experienced barista is at least $295. State any assumption(s) needed for your calculation.

  4. (iv)

    A trainee barista was engaged at the shop. The amount of cheese tea in a cup she made was measured to have mean 498 ml and standard deviation 30 ml. The amount of cheese tea in any cup may be assumed to be independent of the amount of cheese tea in any other cup.

    Find the probability that the mean amount of cheese tea of 40 randomly chosen cups of cheese tea made by the trainee barista does not exceed 500 ml.

Work through it with AI →

It won't just hand you the answer — it asks where you're stuck.How the checker works

Worked solution

13(i)

1. Define the random variables and find the distribution of the total amount of cheese tea in a cup.

Let CC and TT be the amount of cream cheese foam and tea in a cup respectively.
C∼N(72,52)C \sim \mathrm{N}(72, 5^2) and T∼N(421,42)T \sim \mathrm{N}(421, 4^2)

Let X=C+TX = C + T be the total amount of cheese tea in a cup.
E(X)=E(C)+E(T)=72+421=493\mathrm{E}(X) = \mathrm{E}(C) + \mathrm{E}(T) = 72 + 421 = 493
Var(X)=Var(C)+Var(T)=52+42=25+16=41\mathrm{Var}(X) = \mathrm{Var}(C) + \mathrm{Var}(T) = 5^2 + 4^2 = 25 + 16 = 41

Therefore, the distribution is X∼N(493,41)X \sim \mathrm{N}(493, 41).

2. Calculate the probability that the amount of cheese tea is at least 490 ml.

We require P(X≥490)\mathrm{P}(X \ge 490).
Using a graphing calculator,
P(X≥490)=0.68033⋯≈0.680 (to 3 s.f.)\mathrm{P}(X \ge 490) = 0.68033\dots \approx 0.680 \text{ (to 3 s.f.)}

Answer

0.6800.680

13(ii)

1. Define the normal distributions and calculate the probability of a single cup being well-made.

Let XX and YY be the amounts (in ml) of cream cheese foam and tea respectively.
X∼N(72,52)X \sim \mathrm{N}(72, 5^2)
Y∼N(421,42)Y \sim \mathrm{N}(421, 4^2)

Since the amounts are independent, the probability of a cup being well-made is:
P(well-made)=P(65<X<75)×P(416<Y<424)\mathrm{P}(\text{well-made}) = \mathrm{P}(65 < X < 75) \times \mathrm{P}(416 < Y < 424)
=0.64499×0.66772= 0.64499 \times 0.66772
=0.43067 (to 5 s.f.)= 0.43067 \text{ (to 5 s.f.)}

2. Define the binomial distribution and find the probability that more than 5 out of 10 cups are well-made.

Let WW be the number of well-made cups out of 1010.
W∼B(10,0.43067)W \sim \mathrm{B}(10, 0.43067)

P(W>5)=1−P(W≤5)\mathrm{P}(W > 5) = 1 - \mathrm{P}(W \le 5)
=1−0.77381= 1 - 0.77381
=0.22619≈0.226 (to 3 s.f.)= 0.22619 \approx 0.226 \text{ (to 3 s.f.)}

Answer

0.2260.226

13(iii)

1. Formulate the profit for a single cup and find its expected value.

Let CC and TT be the amounts, in ml, of cream cheese foam and tea in one cup respectively.
The profit for one cup, PP, is given by subtracting the cost of the ingredients from the selling price:
P=4.90−(201000)C−(1.201000)T=4.90−0.02C−0.0012TP = 4.90 - \left(\frac{20}{1000}\right)C - \left(\frac{1.20}{1000}\right)T = 4.90 - 0.02C - 0.0012T

The expected profit for one cup is:
E(P)=4.90−0.02E(C)−0.0012E(T)\mathrm{E}(P) = 4.90 - 0.02\mathrm{E}(C) - 0.0012\mathrm{E}(T)
E(P)=4.90−0.02(72)−0.0012(421)=2.9548\mathrm{E}(P) = 4.90 - 0.02(72) - 0.0012(421) = 2.9548

2. Calculate the variance of the profit for a single cup.

Since CC and TT are independent,
Var(P)=(−0.02)2Var(C)+(−0.0012)2Var(T)\mathrm{Var}(P) = (-0.02)^2\mathrm{Var}(C) + (-0.0012)^2\mathrm{Var}(T)
Var(P)=0.0004(52)+0.00000144(42)\mathrm{Var}(P) = 0.0004(5^2) + 0.00000144(4^2)
Var(P)=0.01+0.00002304=0.01002304\mathrm{Var}(P) = 0.01 + 0.00002304 = 0.01002304

3. Define the distribution for the total profit of 100 cups and state the necessary assumption.

Let W=P1+P2+⋯+P100W = P_1 + P_2 + \dots + P_{100} be the total profit for 100 randomly chosen cups.
E(W)=100×2.9548=295.48\mathrm{E}(W) = 100 \times 2.9548 = 295.48
Var(W)=100×0.01002304=1.002304\mathrm{Var}(W) = 100 \times 0.01002304 = 1.002304

Since a linear combination of independent normal variables is also normally distributed,
W∼N(295.48,1.002304)W \sim \mathrm{N}(295.48, 1.002304)

Assumption: The amounts of cream cheese foam and tea from cup to cup (or the profits of each cup) are independent of one another.

4. Calculate the probability that the total profit is at least $295.

We need to find P(W≥295)\mathrm{P}(W \ge 295).
Using a graphing calculator:
P(W≥295)≈0.68419≈0.684 (to 3 s.f.)\mathrm{P}(W \ge 295) \approx 0.68419 \approx 0.684 \text{ (to 3 s.f.)}

Answer

0.6840.684

Assumption: The amounts of ingredients used (or profits) from cup to cup are independent of one another.

13(iv)

1. State the approximate distribution of the sample mean $\bar{X}$ using the Central Limit Theorem.

Let Xˉ\bar{X} be the sample mean amount of cheese tea in 4040 cups.
Since the sample size n=40≥30n = 40 \ge 30 is large, by the Central Limit Theorem, Xˉ\bar{X} is approximately normally distributed.
Xˉ∼N(498,30240)\bar{X} \sim \text{N}\left(498, \frac{30^2}{40}\right)
Xˉ∼N(498,22.5)\bar{X} \sim \text{N}(498, 22.5)

2. Calculate the probability that the sample mean does not exceed $500$ ml.

Using a graphing calculator:
P(Xˉ≤500)=0.66336...\text{P}(\bar{X} \le 500) = 0.66336...
P(Xˉ≤500)≈0.663 (to 3 s.f.)\text{P}(\bar{X} \le 500) \approx 0.663 \text{ (to 3 s.f.)}

Answer

0.6630.663

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1Section B: Statistics [60 marks] A manager claims that the time spent by each customer at his supermarket follows a norm…
  2. 2In this question you should state the parameters of any distributions that you use. Crispy Cream Donut Shop sells 2 type…
  3. 3XX is a normal random variable with mean 1 and variance 1, and YY is also a normal random variable with mean μ\mu and…
  4. 4A credit card company has a special 24-hour telephone hotline for customers to either report a lost credit card or to re…
  5. 5Fishing Company A has two types of fishing vessels, the "Standard" vessel and the "Large" vessel. The amount of fish cau…